# Bell Instability and Cosmic-Ray Acceleration in AGN Ultrafast Outflow Shocks

### v2

@@ 2.2.3 Wind Region
2.2.4 ISM Region
For the ISM region, we assume a stationary homogeneous medium, representative of the narrow line region where relatively low-density gas is observed [^3] AGN outflows including UFOs are suggested to form shocks in the narrow line region (2021PASJ...73.1152J; 2024ApJ...960...41M).. The density and temperature of narrow line region, estimated from emission-line diagnostics, are typically $n_{\text{ISM}}\sim 10^{2}$$10^{4}/\text{cm}^{3}$ and $T_{\text{ISM}}\sim 1.0$$2.5\times 10^{4}~\text{K}$ (1978ApJ...223...56K) (see Sec. 6 of 1997iagn.book.....P). The ISM density is given by
@@ 2.2.4 ISM Region
2.2.5 Initial Magnetic Fluctuations
@@
2.2.6 Injection Rate
The injection rate $\eta$ appearing in Eq. (17) is set to $\eta=10^{-4}$ and $10^{-3}$, as summarized in Tab. 1. Varying $\eta$ changes the number density of CRs, thereby altering the growth efficiency of the NRH instability. Since UFO injection rates are poorly constrained observationally, we adopt $\eta=10^{-4}$$10^{-3}$ as a representative range, guided by studies of DSA in SNRs (2009A&A...499..191T; 2012ApJ...759...12B). This corresponds to cases where a few to tens of percent of the upstream kinetic energy is transferred into CR energy.
Note that $\eta\sim 10^{-3}$ approaches the regime where nonlinear modification of the shock structure and the CR energy spectrum by CR pressure is expected (1997ApJ...485..638M; 1997ApJ...491..584M; 2001RPPh...64..429M; 2005MNRAS.361..907B; 2008MNRAS.385.1946A; 2009ApJ...694..951R; 2013ApJ...775..130S). However, in the present simulations, the MHD momentum and energy equations (Eqs. (3) and (6)) do not yet include CR pressure terms, so these modifications cannot appear. Implementing this back-reaction will be an extension for future work.
2.3 Boundary Conditions
@@
2.4 Numerical Resolution
2.4.1 Spatial Resolution
@@
    The maximum growth wavelength $\lambda_{\text{NRH}}$ of the NRH instability must be resolved with at least 32 cells (see Fig. 21 of 2021ApJ...922....7I).
@@
    The diffusion length $l_{\text{diff}}$ of the lowest-energy CRs ($p_{\text{min}}$) must be resolved with at least 10 cells (see Fig. 20 of 2021ApJ...922....7I).
@@
The maximum growth wavelength $\lambda_{\text{NRH}}$ is given by 2004MNRAS.353..550B as
@@
where $j_{\text{CR}}$ is the CR current at the location where the NRH e-folding number $t_{\text{adv}}/t_{\text{NRH}}$ reaches its maximum. Test calculations estimate this value as $6.2\times 10^{-10}~\text{esu s}^{-1}\text{cm}^{-2}$ for the fiducial model with $(B_{0},~\eta,~\xi_{B,\mathrm{ini}})=(10^{-5}~\mathrm{G},~10^{-4},~0.1)$. The details of the NRH e-folding number are discussed in Sec. 3.1.3.
@@
2.4.2 Momentum-Space Resolution
The momentum space is divided uniformly in logarithmic intervals from $p_{\text{min}}$ to $p_{\text{max}}$. We use $N_{\text{cell},p}=64$ cells, consistent with the convergence tests of 2021ApJ...922....7I.
2.4.3 Time Step Requirement
@@
3 Simulation Results
3.1 Variation of Background Magnetic Field Strength
@@
These results show that for $B_{0}=10^{-5}$ and $10^{-4}~\mathrm{G}$, the maximum energy converges to nearly the same value regardless of $\xi_{B,\mathrm{ini}}$ owing to NRH instability (see Sec. 3.1.3). In contrast, for $B_{0}>10^{-3}~\mathrm{G}$ the NRH instability becomes inefficient, and the maximum energy is determined by $\xi_{B,\mathrm{ini}}$ (see Sec. 3.1.4). This transition corresponds to $\mathcal{M}_{\mathrm{A}}\sim 10^{2}$.
@@
3.1.1 Test-particle prediction from diffusive shock acceleration
In the framework of the DSA model, the energy spectrum of CRs can be analytically derived. In the limit of infinite Mach number, the isotropic distribution function is expressed as (1978MNRAS.182..147B; 1978ApJ...221L..29B; 1983RPPh...46..973D; 1987PhR...154....1B)
@@ 3.1.1 Test-particle prediction from diffusive shock acceleration
3.1.2 Case without NRH instability
@@
3.1.3 Case of weak background magnetic field with NRH instability ($B_{0}=10^{-5},\,10^{-4}~\mathrm{G}$)
@@
Fig. 5 presents the upstream profiles of the CR current density $j_{\text{CR}}$ (left), defined by Eq. (11), and the e-folding number of NRH instability $t_{\text{adv}}/t_{\text{NRH}}$ (right). The e-folding number peaks in a finite region ahead of the shock, and Fig. 4 confirms that magnetic-field amplification occurs predominantly inside this region. The inverse of the linear growth rate of the NRH instability (growth timescale) is represented by (2004MNRAS.353..550B)
@@
Magnetic-field amplification proceeds efficiently in regions where $t_{\text{NRH}}\ll t_{\text{adv}}$. In the case of $B_{0}=10^{-5}~\mathrm{G}$, the e-folding number reaches a maximum of $\sim 5.0$ at $x-x_{\text{sh}}\sim 6.6\times 10^{-1}~\text{pc}$, and significant amplification occurs inside this location [^9] The e-folding number in Fig. 5 (right) represents a local measure of how many times the instability can grow before the shock arrival at each position. The actual number is larger inside the peak because growth initiated farther upstream continues to accumulate over time.. A comparison of the maximum e-folding numbers for weak and strong background magnetic field cases is provided in Tab. 5. The case of strong background fields ($B_{0}=10^{-2}~\mathrm{G}$) is discussed in more depth in Sec. 3.1.4.
@@
In the $B_{0}=10^{-5}~\mathrm{G}$ case, the spectrum at the final time peaks at $\lambda\sim 2.8\times 10^{-4}~\mathrm{pc}$, consistent within a factor of two with the analytically expected NRH maximum growth scale $\lambda_{\text{NRH}}\sim 1.6\times 10^{-4}~\mathrm{pc}$ (Eq. (28)). The final spectrum shows rapid amplification from the shortest scales up to $\lambda_{\text{NRH}}$, followed by a gradual decline toward longer wavelengths. This behavior is explained by the NRH linear growth rate, expressed as (2004MNRAS.353..550B)
@@
3.1.4 Cases with Strong Background Magnetic Fields ($B_{0}=10^{-3},~10^{-2}~\mathrm{G}$)
@@
Even for strong background magnetic fields ($B_{0}=10^{-2}~\mathrm{G}$), the acceleration efficiency agrees with the test-particle DSA prediction (Eq. (37)) within a factor of 2. On the fluid side, however, clear damping of short-wavelength magnetic fluctuations appears, depending on the initial amplitude. As shown in Fig. 8 (left), the case with $\xi_{B,\mathrm{ini}}=0.1$ exhibits strong damping for $\lambda\lesssim 10^{-2}~\mathrm{pc}$, whereas the case with $\xi_{B,\mathrm{ini}}=0.01$ (right) does not. This behavior is consistent with parametric instabilities, particularly stimulated Brillouin scattering, where a parent Alfvén wave decays into a backward Alfvén wave and a sound wave. The growth timescale of stimulated Brillouin scattering is expressed as follows (1978ApJ...224.1013D; 1978ApJ...219..700G; 1993JGR....9819049J; 2024PhRvE.110a5205I),
@@
In the right panel of Fig. 8, the short-wavelength magnetic energy spectra collapse to $P_{\lambda}(\lambda)\simeq\mathrm{const}$, which from Eq. (40) corresponds to $P_{k}(k)\propto k^{-2}$, consistent with both theoretical predictions and solar wind observations. Previous MHD simulations have shown that circularly polarized Alfvén waves with broadband spectra undergo time evolution such that, under magnetically dominated conditions ($c_{\mathrm{s}}\ll v_{\mathrm{A}}$), strong parametric decay into backward Alfvén waves and sound waves occurs, while the instability is suppressed under gas-pressure-dominated conditions ($c_{\mathrm{s}}\gg v_{\mathrm{A}}$) (1992JGR....97.3113U; 1996PhPl....3.4427M; 2001NPGeo...8..159M). Furthermore, 2018JPlPh..84a9006C derived, based on weak turbulence theory, that the scattered Alfvén-wave spectrum scales as $\omega^{-2}$. Observations also report that regions dominated by backward-propagating (sunward) scattered waves exhibit magnetic energy spectra with slopes close to $k^{-2}$ along the background field direction (2008PhRvL.101q5005H; 2009ApJ...698..986P; 2011ApJ...733...76F).
@@ 4 Conclusion
Previous PIC simulations at very low CR maximum energies demonstrated strong magnetic field amplification due to the NRH instability that saturates in nonlinear stage (2014ApJ...783...91C; 2014ApJ...794...46C). By contrast, 2021ApJ...922....7I; 2024ApJ...965..113I showed only moderate amplification in simulations of SNRs consistent with observations. In their simulations, the maximum CR energies were $1~\mathrm{PeV}$ or less. Our study demonstrates that when the maximum energy is even higher ($\lesssim 1~\mathrm{EeV}$), magnetic field amplification due to NRH instability becomes much weaker. Previous PIC simulations have shown that at Alfvén Mach numbers $\mathcal{M}_{\mathrm{A}}\sim 10^{2}$, the magnetic field can be amplified up to $\xi_{B}\sim 60$ (2014ApJ...794...46C). By contrast, in our setup the NRH instability is largely ineffective for $\mathcal{M}_{\mathrm{A}}<10^{2}$ (see Tab. 3). Since the CR current originates from particles escaping upstream near $E_{\max}$, it is natural that the current diminishes with increasing maximum energy, thereby suppressing NRH instability growth. Many earlier works did not explicitly examine the dependence of magnetic amplification on maximum energy, but our results suggest that realistic theoretical models must describe NRH amplification as a function of $E_{\max}$.
@@
    The $\mathrm{p}\gamma$ cooling must remain inefficient (i.e., the AGN photon field must be weak). 2023MNRAS.526..181P showed that EeV-scale acceleration is possible even with $\mathrm{p}\gamma$ cooling if the magnetic energy fraction is high ($\epsilon_{B}\simeq 0.05$). In our simulations, the downstream magnetic energy fraction eventually reaches $\epsilon_{B}^{\mathrm{down}}\simeq 0.04$ for the initial condition $(B_{0},\,\xi_{B,\mathrm{ini}})=(10^{-2}~\mathrm{G},\,0.1)$ (see Tab. 4), similar to their assumed acceleration conditions. However, the maximum CR energy remains sub-EeV (Fig. 11). Thus, achieving EeV energies likely requires a photon field weaker than that in 2023MNRAS.526..181P. The difference is only a factor of a few and is therefore not so serious.
@@
Recent XRISM results indicate mass outflow rates of $\sim 100~M_{\odot}\,\mathrm{yr}^{-1}$ (2025Natur.641.1132X), and both theory and observations suggest that UFOs are clumpy (2013PASJ...65...88T; 2025Natur.641.1132X). Future work must assess how the large mass outflow rates and clumpy structure of UFOs affect efficiency of NRH instability and CR acceleration. Higher mass outflow rates may increase the number of particles contributing to the CR current, thereby enhancing NRH instability. If UFOs are generally clumpy and inhomogeneous, shock geometry would become highly non-uniform, and the acceleration efficiency could differ substantially from the uniform-wind case (e.g., 2012ApJ...744...71I for SNR case). Addressing these issues will require combined theoretical, numerical, and observational efforts.

### v3

@@ II.2.3 Wind Region
II.2.4 ISM Region
For the ISM region, we assume a stationary homogeneous medium, representative of the narrow line region where relatively low-density gas is observed [^3] AGN outflows including UFOs are suggested to form shocks in the narrow line region (Joh et al., 2021; Mizumoto et al., 2024).. The density and temperature of narrow line region, estimated from emission-line diagnostics, are typically $n_{\text{ISM}}\sim 10^{2}$$10^{4}/\text{cm}^{3}$ and $T_{\text{ISM}}\sim 1.0$$2.5\times 10^{4}~\text{K}$ (Koski, 1978) (see Sec. 6 of Peterson (1997)). The ISM density is given by
@@ II.2.4 ISM Region
II.2.5 Initial Magnetic Fluctuations
@@
II.2.6 Injection Rate
The injection rate $\eta$ appearing in Eq. (17) is set to $\eta=10^{-4}$ and $10^{-3}$, as summarized in Tab. 1. Varying $\eta$ changes the number density of CRs, thereby altering the growth efficiency of the NRH instability. Since UFO injection rates are poorly constrained observationally, we adopt $\eta=10^{-4}$$10^{-3}$ as a representative range, guided by studies of DSA in SNRs (Tatischeff, 2009; Berezhko et al., 2012). This corresponds to cases where a few to tens of percent of the upstream kinetic energy is transferred into CR energy.
Note that $\eta\sim 10^{-3}$ approaches the regime where nonlinear modification of the shock structure and the CR energy spectrum by CR pressure is expected (Malkov, 1997a, b; Malkov and Drury, 2001; Blasi et al., 2005b; Amato et al., 2008; Reville et al., 2009; Saito et al., 2013). However, in the present simulations, the MHD momentum and energy equations (Eqs. (3) and (6)) do not yet include CR pressure terms, so these modifications cannot appear. Implementing this back-reaction will be an extension for future work.
II.3 Boundary Conditions
@@
II.4 Numerical Resolution
II.4.1 Spatial Resolution
@@
    The maximum growth wavelength $\lambda_{\text{NRH}}$ of the NRH instability must be resolved with at least 32 cells (see Fig. 21 of Inoue et al. (2021)).
@@
    The diffusion length $l_{\text{diff}}$ of the lowest-energy CRs ($p_{\text{min}}$) must be resolved with at least 10 cells (see Fig. 20 of Inoue et al. (2021)).
@@
The maximum growth wavelength $\lambda_{\text{NRH}}$ is given by Bell (2004) as
@@
where $j_{\text{CR}}$ is the CR current at the location where the NRH e-folding number $t_{\text{adv}}/t_{\text{NRH}}$ reaches its maximum. Test calculations estimate this value as $6.2\times 10^{-10}~\text{esu s}^{-1}\text{cm}^{-2}$ for the fiducial model with $(B_{0},~\eta,~\xi_{B,\mathrm{ini}})=(10^{-5}~\mathrm{G},~10^{-4},~0.1)$. The details of the NRH e-folding number are discussed in Sec. III.1.3.
@@
II.4.2 Momentum-Space Resolution
The momentum space is divided uniformly in logarithmic intervals from $p_{\text{min}}$ to $p_{\text{max}}$. We use $N_{\text{cell},p}=64$ cells, consistent with the convergence tests of Inoue et al. (2021).
II.4.3 Time Step Requirement
@@
III Simulation Results
III.1 Variation of Background Magnetic Field Strength
@@
These results show that for $B_{0}=10^{-5}$ and $10^{-4}~\mathrm{G}$, the maximum energy converges to nearly the same value regardless of $\xi_{B,\mathrm{ini}}$ owing to NRH instability (see Sec. III.1.3). In contrast, for $B_{0}>10^{-3}~\mathrm{G}$ the NRH instability becomes inefficient, and the maximum energy is determined by $\xi_{B,\mathrm{ini}}$ (see Sec. III.1.4). This transition corresponds to $\mathcal{M}_{\mathrm{A}}\sim 10^{2}$.
@@
III.1.1 Test-particle prediction from diffusive shock acceleration
In the framework of the DSA model, the energy spectrum of CRs can be analytically derived. In the limit of infinite Mach number, the isotropic distribution function is expressed as (Bell, 1978; Blandford and Ostriker, 1978; Drury, 1983; Blandford and Eichler, 1987)
@@ III.1.1 Test-particle prediction from diffusive shock acceleration
III.1.2 Case without NRH instability
@@
III.1.3 Case of weak background magnetic field with NRH instability ($B_{0}=10^{-5},\,10^{-4}~\mathrm{G}$)
@@
Fig. 5 presents the upstream profiles of the CR current density $j_{\text{CR}}$ (left), defined by Eq. (11), and the e-folding number of NRH instability $t_{\text{adv}}/t_{\text{NRH}}$ (right). The e-folding number peaks in a finite region ahead of the shock, and Fig. 4 confirms that magnetic-field amplification occurs predominantly inside this region. The inverse of the linear growth rate of the NRH instability (growth timescale) is represented by (Bell, 2004)
@@
Magnetic-field amplification proceeds efficiently in regions where $t_{\text{NRH}}\ll t_{\text{adv}}$. In the case of $B_{0}=10^{-5}~\mathrm{G}$, the e-folding number reaches a maximum of $\sim 5.0$ at $x-x_{\text{sh}}\sim 6.6\times 10^{-1}~\text{pc}$, and significant amplification occurs inside this location [^9] The e-folding number in Fig. 5 (right) represents a local measure of how many times the instability can grow before the shock arrival at each position. The actual number is larger inside the peak because growth initiated farther upstream continues to accumulate over time.. A comparison of the maximum e-folding numbers for weak and strong background magnetic field cases is provided in Tab. 5. The case of strong background fields ($B_{0}=10^{-2}~\mathrm{G}$) is discussed in more depth in Sec. III.1.4.
@@
In the $B_{0}=10^{-5}~\mathrm{G}$ case, the spectrum at the final time peaks at $\lambda\sim 2.8\times 10^{-4}~\mathrm{pc}$, consistent within a factor of two with the analytically expected NRH maximum growth scale $\lambda_{\text{NRH}}\sim 1.6\times 10^{-4}~\mathrm{pc}$ (Eq. (28)). The final spectrum shows rapid amplification from the shortest scales up to $\lambda_{\text{NRH}}$, followed by a gradual decline toward longer wavelengths. This behavior is explained by the NRH linear growth rate, expressed as (Bell, 2004)
@@
III.1.4 Cases with Strong Background Magnetic Fields ($B_{0}=10^{-3},~10^{-2}~\mathrm{G}$)
@@
Even for strong background magnetic fields ($B_{0}=10^{-2}~\mathrm{G}$), the acceleration efficiency agrees with the test-particle DSA prediction (Eq. (37)) within a factor of 2. On the fluid side, however, clear damping of short-wavelength magnetic fluctuations appears, depending on the initial amplitude. As shown in Fig. 8 (left), the case with $\xi_{B,\mathrm{ini}}=0.1$ exhibits strong damping for $\lambda\lesssim 10^{-2}~\mathrm{pc}$, whereas the case with $\xi_{B,\mathrm{ini}}=0.01$ (right) does not. This behavior is consistent with parametric instabilities, particularly stimulated Brillouin scattering, where a parent Alfvén wave decays into a backward Alfvén wave and a sound wave. The growth timescale of stimulated Brillouin scattering is expressed as follows (Derby, 1978; Goldstein, 1978; Jayanti and Hollweg, 1993; Ishizaki and Ioka, 2024),
@@
In the right panel of Fig. 8, the short-wavelength magnetic energy spectra collapse to $P_{\lambda}(\lambda)\simeq\mathrm{const}$, which from Eq. (40) corresponds to $P_{k}(k)\propto k^{-2}$, consistent with both theoretical predictions and solar wind observations. Previous MHD simulations have shown that circularly polarized Alfvén waves with broadband spectra undergo time evolution such that, under magnetically dominated conditions ($c_{\mathrm{s}}\ll v_{\mathrm{A}}$), strong parametric decay into backward Alfvén waves and sound waves occurs, while the instability is suppressed under gas-pressure-dominated conditions ($c_{\mathrm{s}}\gg v_{\mathrm{A}}$) (Umeki and Terasawa, 1992; Malara and Velli, 1996; Malara et al., 2001). Furthermore, Chandran (2018) derived, based on weak turbulence theory, that the scattered Alfvén-wave spectrum scales as $\omega^{-2}$. Observations also report that regions dominated by backward-propagating (sunward) scattered waves exhibit magnetic energy spectra with slopes close to $k^{-2}$ along the background field direction (Horbury et al., 2008; Podesta, 2009; Forman et al., 2011).
@@ IV Conclusion
Previous PIC simulations at very low CR maximum energies demonstrated strong magnetic field amplification due to the NRH instability that saturates in nonlinear stage (Caprioli and Spitkovsky, 2014a, b). By contrast, Inoue et al. (2021, 2024) showed only moderate amplification in simulations of SNRs consistent with observations. In their simulations, the maximum CR energies were $1~\mathrm{PeV}$ or less. Our study demonstrates that when the maximum energy is even higher ($\lesssim 1~\mathrm{EeV}$), magnetic field amplification due to NRH instability becomes much weaker. Previous PIC simulations have shown that at Alfvén Mach numbers $\mathcal{M}_{\mathrm{A}}\sim 10^{2}$, the magnetic field can be amplified up to $\xi_{B}\sim 60$ (Caprioli and Spitkovsky, 2014b). By contrast, in our setup the NRH instability is largely ineffective for $\mathcal{M}_{\mathrm{A}}<10^{2}$ (see Tab. 3). Since the CR current originates from particles escaping upstream near $E_{\max}$, it is natural that the current diminishes with increasing maximum energy, thereby suppressing NRH instability growth. Many earlier works did not explicitly examine the dependence of magnetic amplification on maximum energy, but our results suggest that realistic theoretical models must describe NRH amplification as a function of $E_{\max}$.
@@
    The $\mathrm{p}\gamma$ cooling must remain inefficient (i.e., the AGN photon field must be weak). Peretti et al. (2023) showed that EeV-scale acceleration is possible even with $\mathrm{p}\gamma$ cooling if the magnetic energy fraction is high ($\epsilon_{B}\simeq 0.05$). In our simulations, the downstream magnetic energy fraction eventually reaches $\epsilon_{B}^{\mathrm{down}}\simeq 0.04$ for the initial condition $(B_{0},\,\xi_{B,\mathrm{ini}})=(10^{-2}~\mathrm{G},\,0.1)$ (see Tab. 4), similar to their assumed acceleration conditions. However, the maximum CR energy remains sub-EeV (Fig. 11). Thus, achieving EeV energies likely requires a photon field weaker than that in Peretti et al. (2023). The difference is only a factor of a few and is therefore not so serious.
@@
Recent XRISM results indicate mass outflow rates of $\sim 100~M_{\odot}\,\mathrm{yr}^{-1}$ (Xrism Collaboration et al., 2025), and both theory and observations suggest that UFOs are clumpy (Takeuchi et al., 2013; Xrism Collaboration et al., 2025). Future work must assess how the large mass outflow rates and clumpy structure of UFOs affect efficiency of NRH instability and CR acceleration. Higher mass outflow rates may increase the number of particles contributing to the CR current, thereby enhancing NRH instability. If UFOs are generally clumpy and inhomogeneous, shock geometry would become highly non-uniform, and the acceleration efficiency could differ substantially from the uniform-wind case (e.g., Inoue et al. (2012) for SNR case). Addressing these issues will require combined theoretical, numerical, and observational efforts.